if 48% of the students are female, there are 1,440 female, what is the total number of students in the college?
Answer by ewatrrr(24785)

Hi,
*Note: the 48% tells us what percentage of the TOTAL number of students are female. 48% = 48/100 or .48 written as a decimal
.
Let T represent the TOTAL number fo students
.
The question states the following to be true:
.48*T = 1440
.
Solve for T
T = 3000
.
The total number of students is 3000
.
Check your answer:
.48*3000 = 1440